A farmer drives a tractor from one town to another a distance of 120 kilometers. He drives 10 kilometers per hour faster on the return trip, cutting 1 hour off the time. How fast does he drive each way?
Answer by stanbon(75887)

A farmer drives a tractor from one town to another a distance of 120 kilometers. He drives 10 kilometers per hour faster on the return trip, cutting 1 hour off the time. How fast does he drive each way?
----------------------------------------
1st trip DATA:
distance = 120 km ; rate = x km/hr ; time = d/r = 120/x hrs
--------------------------------
Return trip DATA:
distance = 120 km ; rate = x+10 km/hr ; rime = d/r = 120/(x+10) hrs
-----------------------------------------------------------------
Equation:
1st trip time - return trip time = 1 hr
120/x - 120/(x+10) = 1
120x + 1200 - 120x = x(x+10)
1200 = x^2+10x
x^2 + 10x - 1200 = 0
(x+40)(x-30) = 0
Positive solution:
x = 30 km/hr (rate on the 1st trip)
x+10 = 40 km/hr (rate on the return trip)
============================================
Cheers,
Stan H.